OpenStudy (anonymous):

Medal A space heater is used for 8 h. The heater uses 1600 W per hour. Electricity costs 12.9¢ per kilowatt-hour. Find the cost of using the electric heater. Round to the nearest cent. $ ?

OpenStudy (danjs):

hello

OpenStudy (danjs):

that explains nothing

OpenStudy (anonymous):

i dont think 12.8 * 8 is right

OpenStudy (danjs):

\[\frac{ 12.9 cents }{ 1000 W hr }*\frac{ 1600 W }{ 1 hr }* \frac{ 8 hr }{ 1 }\]

OpenStudy (danjs):

Keep track of your units, if the units do not come out right, in this case

OpenStudy (danjs):

\[\frac{ cost }{ hour }\]

OpenStudy (anonymous):

so how would you multiply this

OpenStudy (danjs):

Just looking at the units, see that you result with Cents / Hour W cancels, and hr cancels

OpenStudy (danjs):

\[\frac{ 12.9 * 1600* 8 }{ 1000 }\frac{ Cents }{ hr }\]

OpenStudy (anonymous):

165120/1000

OpenStudy (danjs):

\[\frac{ 12.9 * 16 * 8 }{ 1 } \frac{ cents }{ hr }\]

OpenStudy (anonymous):

1651.2/1

OpenStudy (danjs):

\[\frac{ 1651.2 cents }{ 1 hr } \]

OpenStudy (danjs):

Remember, the units are more important than the numbers, work through the units first to see if you get the correct resulting unit, in this case cents per hour

OpenStudy (anonymous):

so are the units ok 1651.2 cents / 1 hr

OpenStudy (anonymous):

m i right?

OpenStudy (anonymous):

$1651.20

OpenStudy (danjs):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

It's no wonder you are confused. It's not watts per hour. It's watt times hours, or watt hours. Whoever wrote the problem mislabeled it.